The displacement of a particle S is given by S=A+Bt+Ct^2. Deduce the units of the constants A, B and C.
1
Expert's answer
2019-07-01T09:21:55-0400
"[S]=[A+Bt+Ct^2]"

Thus,


"[A]=[S]=m"

"[B]=\frac{[S]}{[t]}=\frac{m}{s}"

"[C]=\frac{[S]}{[t]^2}=\frac{m}{s^2}"